OpenStudy (anonymous):

what are the x-intercepts of y=(x-2)(x+5)???

OpenStudy (zepp):

Explain to me what is x-intercept? From what you learnt in class.

OpenStudy (hba):

y=o plug it in

OpenStudy (anonymous):

im studying quadratic equations by factoring

OpenStudy (campbell_st):

you need to solve x - 2 = 0 and x + 5 = 0 for the intercepts

OpenStudy (zepp):



OpenStudy (zepp):

x-intercept is all points passing on the x-axis, therefore, y=0

OpenStudy (zepp):

So basically, to find the x-intercepts, you simply replace y by 0 and solve for what values of x, y=0 (or for what values of x, there's points laying on the x-axis.)

OpenStudy (anonymous):

thanks campbell_st!..based off of that i got 2,0 and -5,0 for my final answer

OpenStudy (zepp):

Therefore we can make up two equations: y=(x-2)(x+5) (x-2) = 0 (x+5) = 0 and solve x for each of the equations above.

OpenStudy (anonymous):

thanks zepp!

OpenStudy (anonymous):

x = 2 x = -5 you let y=0 and solve for each x individually.